# Better Not to Propagate: Understanding Edge Uncertainty and Over-smoothing in Signed Graph Neural Networks

### v2

@@ Abstract
Local smoothing, driven by message-passing (MP) under varying levels of homophily, has played a vital role in Graph Neural Networks (GNNs). Recent studies have analyzed this smoothing effect, which reduces node separability after MP, by focusing on the expected distribution of node features. These studies provide theoretical insights into the local smoothing effects of various propagation schemes, including positive, signed, and blocked MPs. Specifically, separability can be theoretically measured using two hyperparameters: a fixed local homophily and a dynamic edge error ratio. However, previous analyses have fixed these values for theoretical justification, which may hinder the optimal selection of MP for GNNs. To address this limitation, we propose a novel method for estimating homophily and edge error ratio, coupled with dynamic selection between blocked and signed propagation during training. Our theoretical analysis, supported by extensive experiments, demonstrates that blocking MP can be more effective than signed propagation under high edge error ratios, thereby improving performance in both homophilic and heterophilic graphs.

### v3

@@ Abstract
Local smoothing, driven by message-passing (MP) under varying homophily levels, is crucial in Graph Neural Networks (GNNs). Recent research has examined this smoothing effect, which diminishes node separability after MP, by analyzing the expected distribution of node features. These studies offer theoretical insights into the local smoothing effects of different propagation methods, such as positive, signed, and blocked MPs. Node separability is theoretically determined by two hyperparameters: a fixed local homophily and a dynamic edge error ratio. However, previous analyses have constrained these values, potentially limiting optimal MP selection for GNNs. To overcome this, we propose a novel method for estimating homophily and edge error ratios, along with a dynamic selection mechanism between blocked and signed propagation during training. Our theoretical analysis with extensive experiments shows that blocking MP outperforms signed propagation under high edge error ratios, improving performance in both homophily and heterophily.